### v1

# EditScore: Unlocking Online RL for Image Editing via High-Fidelity Reward Modeling

Abstract
Instruction-guided image editing has achieved remarkable progress, yet current models still face challenges with complex instructions and often require multiple samples to produce a desired result. Reinforcement Learning (RL) offers a promising solution, but its adoption in image editing has been severely hindered by the lack of a high-fidelity, efficient reward signal. In this work, we present a comprehensive methodology to overcome this barrier, centered on the development of a state-of-the-art, specialized reward model. We first introduce EditReward-Bench, a comprehensive benchmark to systematically evaluate reward models on editing quality. Building on this benchmark, we develop EditScore, a series of reward models (7B–72B) for evaluating the quality of instruction-guided image editing. Through meticulous data curation and filtering, EditScore effectively matches the performance of learning proprietary VLMs. Furthermore, coupled with an effective self-ensemble strategy tailored for the generative nature of EditScore, our largest variant even surpasses GPT-5 in the benchmark. We then demonstrate that a high-fidelity reward model is the key to unlocking online RL for image editing. Our experiments show that, while even the largest open-source VLMs fail to provide an effective learning signal, EditScore enables efficient and robust policy optimization. Applying our framework to a strong base model, OmniGen2, results in a final model that shows a substantial and consistent performance uplift. Overall, this work provides the first systematic path from benchmarking to reward modeling to RL training in image editing, showing that a high-fidelity, domain-specialized reward model is the key to unlocking the full potential of RL in this domain. Github Link: https://github.com/VectorSpaceLab/EditScore.
1 Introduction
Recently, Reinforcement Learning (Li, 2017) has demonstrated immense power in advancing large language models and has also shown remarkable success in the text-to-image (T2I) domain. Works like FlowGRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025) have leveraged RL on flow-matching models to significantly enhance T2I generation across multiple dimensions. Despite these successes, the application of RL to the image editing domain remains largely underexplored. We posit that RL holds significant, untapped potential for image editing. By framing editing as a dynamic trial-and-error process, a policy model can discover novel and more effective editing strategies beyond what is captured in static datasets. Through a well-designed reward mechanism, the model can be progressively steered to better align with user intent and achieve a deeper, more robust editing capability across diverse scenarios.
Despite its theoretical promise, applying online RL to high-resolution image editing remains a formidable and largely unsolved challenge (Wei et al., 2025; Wu et al., 2025a; Ahmadi et al., 2025). The primary obstacle lies in the absence of a suitable reward function: a reliable, efficient, and scalable oracle that can accurately score the quality of an edit given an instruction. A natural consideration would be to employ state-of-the-art, general-purpose Visual Language Models (VLMs) like GPT-5 (OpenAI, 2025) or other proprietary APIs as reward providers. However, this approach is impractical for online RL, with these large VLMs are expensive to query at scale. The second alternative, using powerful open-source VLMs, resolves cost issues. However, our investigations reveal a critical performance gap. As we will demonstrate, even Qwen2.5-VL-72B (Bai et al., 2025) fails to provide a sufficiently accurate and consistent reward signal to effectively guide the policy. Directly using such models as reward functions often causes unstable training or policy collapse, showing that scale alone cannot replace specialized accuracy. Progress thus hinges on an accurate reward model tailored to image editing, yet the community still lacks a powerful open-source option—posing a major barrier to advancing RL-based editing.
In this work, we argue that the key to unlocking online RL for image editing is the development of a specialized, high-fidelity, and efficient reward model. We present a systematic approach to build, validate, and deploy such a model. Our contributions are organized as follows:
First, to enable rigorous and reproducible research, we introduce EditReward-Bench, a comprehensive benchmark for evaluating reward models in image editing. It is organized into four main categories, comprising 13 diverse subtasks ranging from simple attribute edits to complex compositional changes. Each entry is evaluated by expert human raters across three key dimensions: prompt following, consistency, and overall quality. All annotations have passed rigorous inter-annotator agreement checks, establishing a reliable standard for measuring reward model quality.
Second, guided by our benchmark, we develop EditScore, a series of powerful reward models (7B–72B) for evaluating the quality of instruction-guided editing. Through careful data curation and filtering and inference-time scaling, EditScore sets a new state of the art for open-source reward models in this domain, even surpassing the accuracy of leading proprietary VLMs.
To validate its practical utility, we first apply EditScore in a Best-of-$N$ selection experiment across several state-of-the-art editing models, which confirms that our reward model can effectively improve the performance of diverse editors. We further used EditScore for RL training, resulting in a model that shows a substantial and consistent performance uplift over its base model.
In summary, our key contributions are as follows.
  • We propose EditReward-Bench, a new public benchmark for the direct and reliable evaluation of reward models for instruction-guided image editing.
  • We develop and release EditScore, a series of state-of-the-art open-source reward models for instruction-based image editing that, through our self-ensembling strategy, surpasses the accuracy of leading proprietary VLMs on this task.
  • We demonstrate the broad, practical utility of EditScore through a Best-of-$N$ selection experiment, successfully enhancing the performance of multiple diverse base models.
  • We provide a comprehensive analysis of an end-to-end online RL application, identifying key factors for success, including the accuracy and variance of the reward signal.
2 Related Work
Reward Models for Image Generation. Reward models provide signals based on specific preferences and can be applied in various scenarios such as automated evaluation and reinforcement learning. Most research has focused on text-to-image tasks, with works like Wu et al. (2023b); Kirstain et al. (2023); Zhang et al. (2024b) fine-tune CLIP-based model with human preference data and  Ma et al. (2025); Wang et al. (2025b); Gao et al. (2025) employ VLMs as a core component for reward generation. Research on reward models for image editing is relatively underexplored: some works focused on controlled edits based on predefined masks (Ren et al., 2024; Gong et al., 2025). For instruction-guided editing,  Zhang et al. (2024a) fine-tunes the BLIP model (Li et al., 2022) with human annotated reward signals for the image generated only by P2P Hertz et al. (2022) and IP2P Brooks et al. (2023), while Chen et al. (2025) leverages CLIP scores to automatically construct training data for training LLaVA-Next-8B Liu et al. (2024) as reward model. However, they used outdated editing models to generate images that needed to be evaluated, or simply used CLIP to annotate and score the images. Meanwhile, they use limited task categories as the prompt pool. The limitations in the narrow scope of task and generation model coverage make them struggle to support online RL for current models. Furthermore, they are not open-source and cannot be used and evaluated by the public. In contrast, our model enables broader evaluation of editing tasks (13 tasks) while supporting SOTA editing models (e.g., Gemini-2.5-image-preview). Moreover, its generative nature enables it to perform inference-time scaling to enhance scoring accuracy (Liu et al., 2025d).
Image Reward Model Evaluation. Establishing a reliable benchmark is critical as it ensures the objective measurement of reward modeling performance and offers clear guidance for optimization. In previous work, image reward models are evaluated on their accuracy in predicting human preferences for key generative tasks, such as text-to-image generation (Xu et al., 2023; Kirstain et al., 2023; Wu et al., 2023b; a; Ma et al., 2025) and image editing (Ku et al., 2023b; Jiang et al., 2024). Over the past year, generative models such as GPT-Image-1 (Hurst et al., 2024), FLUX-Kontext (Batifol et al., 2025), Qwen-Image-Edit (Wu et al., 2025a) and Nano Banana (Google, 2025) have achieved significant breakthroughs in image editing, particularly in stylization, hybrid edit and text modification. With the growing power of generative models and their expanding editing capabilities, there is an urgent need for reward models that can accurately assess them. This, in turn, necessitates the development of a novel and more comprehensive benchmark for evaluating edit rewards.
3 EditReward-Bench
3.1 Overview
EditReward-Bench is a benchmark specifically designed for systematic evaluation of reward modeling for image editing. It covers 13 diverse editing tasks and spans 11 heterogeneous editing models for data construction, ranging from open-source baselines to state-of-the-art proprietary models (see Table 1 for comparison). The benchmark is distinguished by the following features.
Multi-dimensional Image Evaluation Framework. EditReward-Bench offers three key dimensions for evaluating editing outcomes. These includes Prompt Following (adherence to prompts), consistency (preservation of key visual elements) and overall quality (comparison across all aspects).
Comprehensive Model Performance Spectrum. EditReward-Bench incorporates both state-of-the-art editing models and lower-peforming baselines. It effectively challenges the reward model’s distinguish ability and validates the reward model’s scoring performance on current SOTA models.
Extensive Task Coverage with Real-world Applicability. EditReward-Bench also covers a diverse set of editing tasks that closely align with real-world application scenarios, ensuring authentic and comprehensive evaluation of reward models for image editing.
Table 1: Comparison of EditReward-Bench against existing benchmarks, highlighting its superior scale, data sources, and comprehensive task coverage. The order of subtasks within each tuple under “Task Coverage” corresponds to the order of subtasks listed for each category in Section 3.2.
Feature
ImagenHub
 (Ku et al., 2023b)
GenAI-Bench
 (Jiang et al., 2024)
EditReward-Bench
General Properties
Size 2,864 919 3,072
Multi-Dimensional Eval.
Proprietary Model Data
Task Coverage (Conceptual Groups)
[…]
3.2 Construction Protocol
To create a robust and comprehensive benchmark, we focused on three key pillars: the diversity of editing tasks, the variety of editing models, and the granularity of our evaluation criteria.
First, to ensure comprehensive task coverage, we structure our benchmark into four main categories, comprising 13 distinct subtasks curated from established datasets like GEdit-Bench-EN (Liu et al., 2025b) and ImgEdit-Bench (Ye et al., 2025). These categories are designed to span a wide spectrum of complexity. They are: 1) Subject, which includes fundamental tasks like subject addition, subject removal and subject replace; 2) Appearance, covering shape-preserving edits such as color alteration, material modification, style transfer and tone transformation; 3) Scene, which tests understanding of image layout through tasks like background change and extract; and 4) Advanced, the most challenging category, which requires advanced reasoning for tasks like portrait beautification, text modification, motion change and hybrid edit.
Second, to populate our benchmark with a diverse distribution of edited images, the candidate pool of outputs were generated by a diverse array of 11 generative models for data construction: Step1X-Edit (Liu et al., 2025b), Step1X-Edit v1.1 (Liu et al., 2025b), Qwen-Image-Edit (Wu et al., 2025a), OmniGen2 (Wu et al., 2025b), FLUX-Kontext-dev (Batifol et al., 2025), FLUX-Kontext-pro (Batifol et al., 2025), Bagel (Deng et al., 2025), MagicBrush (Zhang et al., 2023), Omnigen (Xiao et al., 2025), gpt-image-1 (Hurst et al., 2024), and Gemini-2.5-image-preview (Google, 2025), including both open-source and state-of-the-art proprietary editors.
Finally, recognizing that the quality of editing results is not monolithic, we designed a multi-dimensional evaluation scheme for better interpretability as inspired by (Ku et al., 2023a; Liu et al., 2025b). Each editing result is assessed along three distinct axes: Prompt Following (PF): measuring how faithfully the edit executes the given instruction. Consistency (C): assessing the preservation of unedited image regions. Overall Quality (O): providing a holistic comparison of edits by accounting for all relevant aspects.
3.3 Annotation Pipeline
To ensure the reliability of EditReward-Bench, we designed a rigorous human annotation pipeline. The process was conducted exclusively by experts in generative AI to guarantee data quality and consistency, as illustrated in Figure 1.
Before the formal annotation, all raters scored a small subset of the entire sample pool to discuss disagreements and collectively established unified scoring guidelines. Then the raters proceeded to the formal annotation of the remaining samples. For each input with five output images randomly sampled from the candidate pool, they were asked to rank them. Our protocol uniquely allowed them to group outputs of similar quality into the same tiers; for instance, a ranking of $3|12|45$ places output 3 in the top tier, outputs 1 and 2 tied in the second, and 4 and 5 tied in the third. Crucially, each data point was annotated by two raters, and only entries where their rankings were in complete agreement were accepted into the final benchmark, ensuring maximum reliability.
These tiered rankings were then decomposed into pairwise preference tuples. Following the $3|12|45$ example, this conversion yields preference pairs where a higher-tiered item is preferred over a lower-tiered one, such as $(3,1)$, $(3,2)$, $(1,4)$ and $(2,5)$. This method correctly excludes comparisons between items within the same tier (e.g., $(1,2)$), as they represent equal quality. This process resulted in a large-scale, high-fidelity dataset of pairwise comparisons, comprising 944 pairs for prompt following, 890 for consistency, and 1,238 for overall quality, for a total of 3,072 preference pairs. A detailed overview of our custom-built annotation interface is provided in Appendix F.
Figure 1: Illustration of the annotation process. Annotators are presented with five candidate output images and are asked to rank them according to three evaluation dimensions. The final ranking is determined through consensus among multiple annotators. For example, $1|234|5$ indicates that the first image is preferred over images 2, 3 and 4, which are in turn preferred over image 5.
4 Method
4.1 Edit Reward
4.1.1 The Approach to Reward Modeling
Inspired by (Wang et al., 2025b; a; Wu et al., 2025c), we formulate reward modeling as a conditional textual generation task. We fine-tune models from the powerful Qwen2.5-VL series (Bai et al., 2025) on a standard autoregressive objective to act as specialized evaluators. Our model EditScore takes (Instruction, Input Image, Output Image) as input, (Reasoning, Scalar Score) as output. The chain-of-thought (Wei et al., 2022) process of output not only enhances the model’s interpretability but also demonstrably improves the accuracy of the final scores.
To ensure EditScore can perform a comprehensive evaluation of editing tasks, we adopt the VIEScore framework (Ku et al., 2023a), prompting the model in parallel to assess two orthogonal aspects. The first is Semantic Consistency (SC), which evaluates the degree of instruction following and consistency, ensuring specified objects are correctly modified while unmentioned regions remain preserved. The second is Perceptual Quality (PQ), which assesses overall image quality, focusing on photorealism and the absence of artifacts. The final score is the geometric mean of the SC and PQ scores ($S_{final}=\sqrt{S_{SC}\cdot S_{PQ}}$), yielding a balanced and granular reward signal.
4.1.2 Inference-time ensembling Strategy
Recent work (Snell et al., 2024; Liu et al., 2025c) has shown that the performance of LLMs can be substantially improved by increasing the computational budget at test time. Building on this insight, we introduce an inference-time ensembling strategy for our generative reward model to systematically boost its evaluation accuracy and enhance reward stability. Our approach is straightforward: for a given input triplet $\mathbf{z}=(\text{Instruction, Input Image, Output Image})$, we perform $K$ independent, stochastic forward passes through EditScore, and each pass $i$ generates a pair $(\text{reasoning}_{i},s_{i})$. Then we aggregate only the scalar scores $\{s_{1},s_{2},\dots,s_{K}\}$ to compute the final ensembled score, $S_{final}$. An intuitive explanation (Liu et al., 2025c) for this effectiveness is that each of the $K$ generated reasonings can be viewed as distinct judgment perspectives. Aggregating the scores derived from these diverse perspectives allows the final reward to more accurately reflect the true quality of an edit, leading to significant scaling effectiveness.
4.2 Data Construction Pipeline
The training data for reward model and subsequent reinforcement learning are constructed following the similar procedure. For reinforcement learning, the training data only requires the input images and instructions, whereas training the reward model additionally requires the generated outputs paired with their corresponding rewards. The construction process involves three steps as follows.
Step I: Images selection and instructions creation. We start by selecting and filtering a diverse set of high-quality images as editing inputs. Next, we construct a series of reference instructions for various editing tasks. Using the input images and random selected reference instructions as guidance, we prompt Qwen-2.5-VL-72B to generate editing instructions that are task-consistent with the references. After generating a large set of image-instruction pairs, we applying a K-center greedy algorithm (Sener & Savarese, 2018) to select 1000 semantically diverse samples per task for reward generation.
Finally, we constructed 70,000 data samples for training the reward model and 60,000 samples for reinforcement learning training. The data for the reward model requires further annotation as follows.
Step II: Candidates output generation. We generate candidates output using 5 distinct editing models random selected from our model pool.
Step III: Annotate and filtering. For each candidate output, we use GPT-4.1 to annotate scores. Following VIEScore (Ku et al., 2023a), we score each output on SC and PQ and provide reasons for the assigned scores. Next, we apply filtering to the reward samples. It is conducted from two perspectives: (i) filtering based on group-wise maximum scores to remove unachievable editing tasks, and (ii) filtering by group standard deviation to remove cases with low discriminability.
5 Reward Model Performance on EditReward-Bench
5.1 Experimental Setup
Our final EditScore model is obtained by fine-tuning Qwen2.5-VL with LoRA (Hu et al., 2022) on our curated dataset (see Section 4.2). For evaluation, we adopt the VIEScore (Ku et al., 2024) prompt template with two key modifications: (i) enforcing a reasoning-before-scoring format, (ii) expanding the score range from $[0,10]$ to $[0,25]$. Following the formulation of VIEScore, we derive Prompt Following ($S_{PF}$) and Consistency ($S_{C}$) from its Semantic Consistency metric $S_{SC}$, while Overall Quality ($O$) is directly taken from the final score $S_{final}$. The detailed prompt templates are shown in Appendix H.
Table 2: Benchmark results on EditReward-Bench, reporting both overall pairwise accuracy and a fine-grained breakdown across four categories of edit capabilities. Notably, EditScore achieves superior performance even with its compact 7B size. Avg@4 denotes the average score over 4 forward passes.
Model GPT-4.1 GPT-5 Gemini-2.5 Qwen2.5-VL EditScore-7B EditScore-32B EditScore-72B
Metric Pro 7B 32B 72B Base Avg@4 Base Avg@4 Base Avg@4
Overall PF 0.673 0.777 0.703 0.458 0.498 0.540 0.592 0.722 0.638 0.736 0.635 0.755
C 0.602 0.669 0.560 0.325 0.376 0.435 0.591 0.720 0.556 0.704 0.586 0.735
O 0.705 0.755 0.722 0.432 0.563 0.621 0.659 0.727 0.680 0.733 0.703 0.763
Subject PF 0.615 0.707 0.712 0.414 0.527 0.523 0.590 0.691 0.625 0.703 0.612 0.708
C 0.520 0.538 0.465 0.317 0.414 0.394 0.585 0.666 0.473 0.627 0.524 0.639
O 0.679 0.708 0.765 0.460 0.570 0.594 0.740 0.771 0.703 0.754 0.721 0.807
Appear. PF 0.673 0.762 0.631 0.422 0.393 0.390 0.573 0.682 0.587 0.714 0.584 0.733
C 0.668 0.714 0.577 0.335 0.320 0.416 0.612 0.730 0.591 0.764 0.623 0.778
O 0.709 0.756 0.700 0.470 0.514 0.559 0.663 0.714 0.669 0.710 0.697 0.736
Scene PF 0.763 0.852 0.766 0.433 0.611 0.690 0.744 0.821 0.789 0.870 0.788 0.908
C 0.682 0.741 0.675 0.236 0.482 0.429 0.735 0.835 0.627 0.787 0.695 0.797
O 0.773 0.841 0.693 0.357 0.673 0.713 0.789 0.774 0.764 0.816 0.794 0.837
Advanced PF 0.673 0.806 0.736 0.541 0.524 0.627 0.536 0.736 0.625 0.717 0.625 0.734
C 0.556 0.687 0.557 0.367 0.351 0.488 0.503 0.693 0.548 0.658 0.541 0.733
O 0.686 0.746 0.724 0.410 0.553 0.657 0.529 0.683 0.631 0.699 0.650 0.721
5.2 Main Results
Results summarized in Table 2 reveal a significant performance gap between proprietary models and open-source counterparts. Leading proprietary models such as GPT-4.1, GPT-5 and Gemini-2.5-Pro form a distinct upper tier, achieving pairwise accuracies in the 0.7-0.75 range across all dimensions. This confirms their strong, albeit imperfect, zero-shot capabilities for this nuanced task. In particular, we find that VLMs are generally stronger at assessing Prompt Following (PF) than Consistency (C), as the latter requires fine-grained comparisons between the input and output images.
In stark contrast, even the largest and most capable open-source models exhibit notable limitations. The Qwen2.5-VL series shows a clear scaling trend, yet the 72B-parameter variant still falls short of 0.612 overall accuracy and performs worse than random chance in Consistency judgment. The smaller 7B and 32B models fare even worse, underscoring the inadequacy of off-the-shelf open-source VLMs as reliable reward signals for fine-grained editing tasks. By contrast, our EditScore achieves substantial improvements: the 7B variant surpasses the 10x larger Qwen2.5-VL-72B, while the 72B variant matching the score of GPT-4.1. Moreover, scaling inference-time compute with self-ensemble (Avg@4) further boosts performance across all model sizes, with EditScore-72B establishing the state of the art on EditReward-Bench.
5.3 Effective Inference-time scaling of EditScore
Figure 2: Self-ensembling offers a superior efficiency-performance trade-off compared to simply scaling model parameters. The colored solid lines show the performance scaling of our 7B, 32B, and 72B models as the number of ensemble passes (K) increases. The gray dashed line connects the single-pass (K=1) performance of these models, serving as a baseline for scaling model size alone. The results clearly indicate that scaling the number of forward passes yields a significantly higher accuracy gain per unit of computational cost.
As shown in Table 2, both scaling model size and inference-time compute improve the performance of EditScore. To further investigate which dimension contributes more effectively, we normalize the comparison by using FLOPs as a proxy for inference cost. Figure 2 presents the scaling trends of EditScore across different model sizes: colored solid lines denote inference-time compute scaling, while gray dotted lines indicate parameter scaling. We observe that, across all model sizes, scaling inference-time compute yields greater performance gains than scaling parameters. Moreover, in practical deployment, the self-ensemble strategy allows the profiling stage to be shared within the ensemble group, which improving efficiency further.
5.4 Ablation Study
We study two key factors in EditScore design: score range granularity and output format (Table 3). Increasing the target score range generally improves pairwise accuracy for both GPT-4.1 and GPT-5, peaking around $[0,25]$, while overly large ranges hurt performance due to regression difficulty. In parallel, requiring the model to generate a rationale before the numeric score (“reasoning + score”) consistently outperforms direct scoring, yielding a +0.038 accuracy gain for EditScore-7B (0.621 → 0.659). These findings highlight that both an appropriately chosen score range and reasoning-first output are crucial for maximizing accuracy.
Table 3: The effect of score range granularity and reason first on the pairwise accuracy of VLMs on EditReward-Bench. The ranges ($[0,x]$) indicate the target score space for model outputs. The lightly shaded column highlights our chosen configuration for optimal performance.
Method Score Range Output Format
$[0,10]$ $0,20] \columncolorlightgray[$0,25][0,25] $0,30] \columncolorlightgrayreasoning$ + score score
[…]
6 Application of EditScore in Image Editing
6.1 Experimental Setup
Evaluation Method. To evaluate the effectiveness of EditScore in improving image editing models, we design two experiments. (1) best-of-$N$ selection: EditScore is used as a selector, where the editing model generates multiple candidate outputs per input and EditScore chooses the best one. We evaluate the gain on three popular models: OmniGen2 (Wu et al., 2025b), Flux.1-Kontext-dev Batifol et al. (2025) and Qwen-Image-Edit Wu et al. (2025a). (2) Reinforcement learning: EditScore is employed directly as a reward model to fine-tune OmniGen2 via an additional RL stage, demonstrating its utility as a training signal. We adopt two widely used image editing benchmarks—GEdit-Bench (Liu et al., 2025b) and ImgEdit-Bench (Ye et al., 2025)—which cover a diverse range of practical editing tasks, to assess the improvements brought by EditScore. For efficiency, we use 7B variant of EditScore for experiments with optional self-ensemble.
RL Training. OmniGen2 Wu et al. (2025b) is chosen as the base model due to its strong potential for further improvement, as evidenced in the best-of-$N$ experiments (Section 6.2). We employ Flow-GRPO Liu et al. (2025a) with the following hyperparameters: sampling steps $T=20$, group size $G=12$, number of unique prompts = 24, noise level $\sigma=0.9$, and KL weight $\beta=0.04$. We reformulate the standard Flow-GRPO equation to align with OmniGen2’s definition; detailed derivations are provided in Appendix B and C.
6.2 Validating EditScore Utility via best-of-$N$ Section
Figure 3: EditScore as a superior reward signal for image editing. (a) Using EditScore to select the best sample among multiple outputs effectively improves VIEScore, with OmniGen2 showing the largest gain. (b) Incorporating EditScore into RL training yields stable and significant performance improvements, even surpassing the much larger Qwen2.5-VL-72B. (c) RL training benefits from self-ensembling, which enhances the evaluation accuracy of EditScore across diverse settings.
Before implementing full RL training, we first validate the utility of EditScore in a controlled setting. In this experiment, the editing model generates $N$ candidate outputs per input of GEdit-Bench, and EditScore selects the best one. This setup directly measures the reward model’s ability to identify high-quality edits. Figure 3(a) reports benchmark performance as a function of $N$. We evaluate three base models using both single-pass EditScore (solid lines) and the stronger Avg@4 version (dashed lines). The results highlight three key findings: (1) stronger reward models consistently yield better selections, (2) performance gains vary by base model, with Qwen-Image-Edit showing the least improvement due to prior post-training stage (Wu et al., 2025a), and (3) OmniGen2 exhibits the largest absolute potential for improvement, motivating its choice as the base model for subsequent RL experiments.
6.3 Main RL Results and Analysis
Having established the utility of EditScore in a controlled best-of-$N$ setting and identified OmniGen2 as a high-potential candidate, we now proceed to the core of our investigation: leveraging EditScore as a direct reward signal for online RL fine-tuning.
6.3.1 Unlocking Stable RL with a High-Fidelity Reward Signal
Our experiments reveal that successful RL fine-tuning in the image editing domain is not a given; it hinges critically on the quality of the reward signal.
The Decisive Role of Reward Model Fidelity. The central thesis of our work is that a high-fidelity, specialized reward model is a prerequisite for effective RL. Figure 3(b) provides a stark validation of this claim. When attempting to use general-purpose, off-the-shelf VLMs (Qwen2.5-VL series) as reward providers, the training process is highly unstable and the policy model fails to converge, even with the 72B parameter model. This failure highlights a crucial insight: raw VLM capabilities do not readily translate into a usable reward signal. In stark contrast, our distilled EditScore provides a clean, consistent, and informative reward, enabling a stable and highly effective learning trajectory.
Unlocking Further Gains with Self-Ensembling. Beyond the model itself, the inference-time strategy for deploying the reward model is also impactful. Figure 3(c) compares the learning curves using a single-pass EditScore versus our 4-pass self-ensemble strategy. The ensembled reward signal demonstrably leads to faster convergence and a superior final policy performance. This confirms a vital connection: improvements in the reward model’s standalone accuracy (as shown in our main results) directly translate to more efficient and effective downstream RL training.
6.4 Analysis of Annotation Sources and Reward Variance
Table 4: Comparison of reward models based on annotator source. We analyze reward models (RMs) trained on data from two different annotators: GPT-4.1 and GPT-5. The resulting policy’s performance is evaluated by both annotators on GEdit-Bench (Liu et al., 2025b).
Annotator Component Reward Model Metrics Policy Performance $\uparrow$
Acc. (O) $\uparrow$ Score Std. GPT-4.1 Score GPT-5 Score
GPT-4.1 (Better for RL) Annotator (Itself) 0.705 3.309
$\hookrightarrow$ EditScore (RM) 0.637 2.868
   $\hookrightarrow$ OmniGen2 (Policy) 6.375 5.834
GPT-5 (Worse for RL) Annotator (Itself) 0.755 2.942
$\hookrightarrow$ EditScore (RM) 0.638 2.533
   $\hookrightarrow$ OmniGen2 (Policy) 6.292 5.768
To understand the role of annotators in shaping RL performance, we trained two distinct EditScore-7B reward models on identical data subsets, annotated by GPT-4.1 and GPT-5 separately. While GPT-5 offers superior annotation accuracy, the higher score variance of GPT-4.1 provides a more effective learning signal. As shown in Table 4, EditScore trained on GPT-4.1’s labels, despite a marginal dip in accuracy (0.637 vs. 0.638), inherits a substantially higher score standard deviation (2.868 vs. 2.533). This high-variance reward signal leads to a demonstrably better policy after RL fine-tuning. Crucially, this improvement is consistent between both GPT-4.1 (6.375 vs. 6.292) and GPT-5 (5.834 vs. 5.768) evaluations, excluding potential evaluation bias from GEdit-Bench (Ye et al., 2025). Our results thus uncover a key insight: reward signal variance, rather than absolute annotator accuracy, can be the dominant factor for successful optimization of the RL-based model, corroborating similar findings from Razin et al. (2025).
6.4.1 Quantifying the Final Performance Gains
Table 5: Performance comparison of models with RL enhancement, where our EditScore provides substantial gains for OmniGen2 across two widely adopted benchmarks.
Model GEdit-Bench-EN ImgEdit-Bench
SC PQ O O
UniPic2-SD3.5M-Kontext 6.31 3.95
UniPic2-SD3.5M-Kontext RL w/ GPT-4.1 6.55 (+0.24) 4.04 (+0.09 )
OmniGen2 6.72 7.20 6.28 3.40
OmniGen2 RL w/ EditScore 7.20 (+0.48) 7.46 (+0.26) 6.68 (+0.40) 3.63 (+0.23)
Table 5 summarizes the substantial gains achieved by OmniGen2 after RL fine-tuning with EditScore. We also provide a visualization of the results before and after RL in the Appendix 5.
On GEdit-Bench-EN, RL raises the overall score from 6.28 to 6.68, a notable improvement of +0.40, with particularly strong gains in Semantic Control (+0.48). On the more challenging ImgEdit-Bench, we also observe a meaningful increase of +0.23. For comparison, we include UniPic2 with EditReward, showing that the improvements achieved by EditScore are both competitive and substantial. These results provide strong evidence that a high-fidelity reward model such as EditScore is essential for unlocking the full potential of online RL in image editing.
7 Conclusion
In this work, we addressed the critical bottleneck for RL in image editing: the lack of a reliable reward signal. We established a comprehensive benchmark for reward model evaluation. Guided by our benchmark, we developed EditScore, a family of specialized generative reward models that deliver a robust, high-fidelity signal. Crucially, EditScore is not only efficient for Best-of-$N$ selection but is the key to unlocking stable online RL where previous open-source models fail. By releasing both EditReward-Bench and EditScore, we provide a foundational toolkit for future research into RL-based image editing and more nuanced reward modeling.
Appendix A Statement on the Use of Large Language Models (LLMs)
In adherence to the ICLR 2026 submission guidelines, this section details the use of a Large Language Model (LLM) assistant during the preparation of this manuscript. The LLM, acting as a research and writing co-pilot, played a significant role in refining the manuscript’s structure, language, and presentation. The authors maintained full intellectual control throughout the process and take complete responsibility for all content.
The precise role of the LLM can be categorized as follows:
  • Manuscript Writing and Polishing: The authors wrote the initial drafts for all sections of the paper, providing the key technical details, experimental results, and core arguments. The LLM was then used extensively as an interactive writing assistant to:
    • Enhance Clarity and Conciseness: Rephrasing long or complex sentences to improve readability and flow.
    • Improve Academic Tone: Suggesting more formal and professional vocabulary and sentence structures appropriate for a top-tier conference submission.
    • Correct Grammar and Syntax: Performing comprehensive proofreading to identify and correct grammatical errors, typos, and awkward phrasing.
    • Suggest Alternative Phrasing: Providing multiple options for expressing a single idea to avoid repetitive language).
  • Technical Formalization: For complex mathematical sections, such as the derivation of the SDE from the ODE in Appendix B, the authors provided the core mathematical steps and handwritten notes. The LLM assisted in translating these steps into a clear, well-structured narrative and formatting them professionally in LaTeX. The LLM did not generate novel mathematical proofs but rather helped in their presentation.
Throughout this collaborative process, every suggestion and piece of text generated by the LLM was critically reviewed, edited, and approved by the human authors. The authors are solely responsible for the scientific validity, originality, and all claims made in this paper. The LLM is not considered an author.
Appendix B Derivation of the SDE Formulation for Flow Matching
This appendix derives the stochastic differential equation (SDE) underlying policy sampling in our Flow Matching model. Unlike prior work such as Flow-GRPO (Liu et al., 2025a), our base model OmniGen2 defines the probability path from noise at $t=0$ ($\mathbf{x}_{0}$) to data at $t=1$ ($\mathbf{x}_{1}$), reversing the convention in some other studies. This leads to a different drift term in the resulting SDE. The derivation proceeds in three steps: (i) connecting the deterministic ODE and a general SDE via the Fokker–Planck equation, (ii) relating the score term to the learned vector field, and (iii) constructing the discretized transition kernel.
B.1 Connecting the Probability Flow ODE to an SDE
Our Flow Matching model  (Lipman et al., 2023) is trained to approximate the vector field $v_{t}(\mathbf{x}_{t})$ of a probability flow ODE, which deterministically transports samples from a noise distribution $p_{0}$ to a data distribution $p_{1}$:
$d\mathbf{x}_{t}=v_{t}(\mathbf{x}_{t})dt.$ (1)
We wish to find an equivalent SDE of the general form
$d\mathbf{x}_{t}=f(\mathbf{x}_{t},t)dt+G(t)dW_{t},$ (2)
that generates the same marginal probability density path $p_{t}(\mathbf{x})$ for all $t\in[0,1]$. Here, $f$ is the drift term, $G(t)$ is the diffusion coefficient (we assume it is state-independent), and $W_{t}$ is a standard Wiener process.
The evolution of the probability density $p_{t}(\mathbf{x})$ is described by the Fokker-Planck equation (Song et al., 2021). For the deterministic ODE in equation 1, it is:
$\frac{\partial p_{t}(\mathbf{x})}{\partial t}=-\nabla\cdot\left[v_{t}(\mathbf{x})p_{t}(\mathbf{x})\right].$ (3)
For the SDE in equation 2, the Fokker-Planck equation is:
$\frac{\partial p_{t}(\mathbf{x})}{\partial t}=-\nabla\cdot\left[f(\mathbf{x},t)p_{t}(\mathbf{x})\right]+\frac{1}{2}\nabla^{2}\left[G(t)^{2}p_{t}(\mathbf{x})\right].$ (4)
For the two processes to be equivalent (i.e., to share the same $p_{t}(\mathbf{x})$), the right-hand sides of equation 3 and equation 4 must be equal. This allows us to solve for the SDE drift term $f(\mathbf{x},t)$:
$-\nabla\cdot\left[v_{t}p_{t}\right]$ $=-\nabla\cdot\left[fp_{t}\right]+\frac{1}{2}\nabla\cdot\left[\nabla(G^{2}p_{t})\right]$ (5)
$\nabla\cdot\left[v_{t}p_{t}\right]$ $=\nabla\cdot\left[fp_{t}-\frac{1}{2}G^{2}\nabla p_{t}\right]$ (6)
$v_{t}p_{t}$ $=fp_{t}-\frac{1}{2}G^{2}\nabla p_{t}.$ (7)
Dividing by $p_{t}$ and rearranging yields the expression for the drift:
$f(\mathbf{x}_{t},t)=v_{t}(\mathbf{x}_{t})+\frac{G(t)^{2}}{2}\nabla\log p_{t}(\mathbf{x}_{t}).$ (8)
Substituting this back into equation 2, we obtain the equivalent SDE:
$d\mathbf{x}_{t}=\left[v_{t}(\mathbf{x}_{t})+\frac{G(t)^{2}}{2}\nabla\log p_{t}(\mathbf{x}_{t})\right]dt+G(t)dW_{t}.$ (9)
We define that noise at $t=0$ ($\mathbf{x}_{0}$) and ending with the data sample at $t=1$ ($\mathbf{x}_{1}$), so equation 9 don’t need a reverse SDE. This SDE’s drift consists of the original ODE vector field plus a term proportional to the score function, $\nabla\log p_{t}(\mathbf{x}_{t})$.
B.2 Expressing the Score via the Learned Vector Field
The SDE in equation 9 is not yet practical for sampling, as the score function $\nabla\log p_{t}(\mathbf{x}_{t})$ is unknown. However, for the specific probability path used in Flow Matching, this score can be expressed entirely in terms of the known vector field $v_{t}(\mathbf{x}_{t})$, which is approximated by our trained model $v_{\theta}$.The derivation follows a similar approach to that for stochastic interpolants (Albergo et al., 2023).
Our derivation begins with the linear interpolation path: $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$, where $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\mathbf{x}_{1}$ is a data sample. The velocity of this path is defined as the conditional expectation of the instantaneous change:
$v_{t}(\mathbf{x}_{t})=\mathbb{E}\left[\frac{d\mathbf{x}_{t}}{dt}\Big|\mathbf{x}_{t}\right].$ (10)
Since the time derivative is $\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-\mathbf{x}_{0}$, the vector field becomes:
$v_{t}(\mathbf{x}_{t})=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t}]-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}].$ (11)
A key identity, derived from Tweedie’s formula in score-based modeling, connects the marginal score $\nabla\log p_{t}(\mathbf{x}_{t})$ to the posterior mean of the initial noise sample $\mathbf{x}_{0}$:
$\nabla\log p_{t}(\mathbf{x}_{t})=-\frac{\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{1-t}.$ (12)
Our goal is thus to express $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ using $v_{t}(\mathbf{x}_{t})$. To do this, we take the conditional expectation of the path definition itself:
$\mathbb{E}[\mathbf{x}_{t}|\mathbf{x}_{t}]=\mathbf{x}_{t}=(1-t)\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]+t\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t}].$ (13)
We now have a system of two linear equations ( equation 11 and equation 13) with two unknowns ($\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ and $\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t}]$). We can solve this system for $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$. Rearranging equation 11 gives $\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t}]=v_{t}(\mathbf{x}_{t})+\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$. Substituting this into equation 13:
$\mathbf{x}_{t}$ $=(1-t)\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]+t(v_{t}(\mathbf{x}_{t})+\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$
$\mathbf{x}_{t}$ $=(1-t+t)\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]+tv_{t}(\mathbf{x}_{t})$
$\mathbf{x}_{t}$ $=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]+tv_{t}(\mathbf{x}_{t}).$ (14)
This gives us the desired expression:
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{t}-tv_{t}(\mathbf{x}_{t}).$ (15)
Finally, substituting equation 15 back into the score identity ( equation 12), we arrive at the final, practical expression for the score function:
$\nabla\log p_{t}(\mathbf{x}_{t})=-\frac{\mathbf{x}_{t}-tv_{t}(\mathbf{x}_{t})}{1-t}.$ (16)
This result allows us to compute the score at any point $(\mathbf{x}_{t},t)$ using only the output of our trained vector field model $v_{\theta}(\mathbf{x}_{t},t)$.
B.3 Final SDE and Discretized Transition Probability
We now substitute the practical expression for the score equation 16 into the SDE from equation 9. For simplicity, let’s assume a constant diffusion $G(t)=\sigma$:
$d\mathbf{x}_{t}=\left[v_{t}(\mathbf{x}_{t})-\frac{\sigma^{2}}{2}\frac{\mathbf{x}_{t}-tv_{t}(\mathbf{x}_{t})}{1-t}\right]dt+\sigma dW_{t}.$ (17)
This is our final SDE, where the drift is expressed entirely in terms of the learned vector field $v_{t}$ (approximated by $v_{\theta}$). Unlike some prior work that requires deriving a reverse-time SDE, we can directly use this forward-time SDE for model sampling.
To implement this for our RL policy, we discretize equation 17 using the Euler-Maruyama scheme with a step size $\Delta t$:
$\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\left[v_{\theta}(\mathbf{x}_{t},t)-\frac{\sigma^{2}}{2}\frac{\mathbf{x}_{t}-tv_{\theta}(\mathbf{x}_{t},t)}{1-t}\right]\Delta t+\sigma\sqrt{\Delta t}\cdot\bm{\epsilon},$ (18)
where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
This defines the stochastic transition probability of our policy, $\pi_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},\mathbf{c})$. It is a Gaussian distribution:
$\pi_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},\mathbf{c})=\mathcal{N}(\mathbf{x}_{t+\Delta t};\bm{\mu}(\mathbf{x}_{t},t),\Sigma),$ (19)
with mean and covariance given by:
$\bm{\mu}(\mathbf{x}_{t},t)$ $=\mathbf{x}_{t}+\left[v_{\theta}(\mathbf{x}_{t},t)-\frac{\sigma^{2}}{2}\frac{\mathbf{x}_{t}-tv_{\theta}(\mathbf{x}_{t},t)}{1-t}\right]\Delta t$ (20)
$\Sigma$ $=\sigma^{2}\Delta t\cdot\mathbf{I}.$ (21)
This completes the derivation from the deterministic ODE to a concrete, sampleable stochastic policy for reinforcement learning.
Appendix C GRPO on SDE Flow Matching
With the stochastic policy $\pi_{\theta}$ defined by the SDE in Appendix B, we optimize our generative model using GRPO (Shao et al., 2024). GRPO is efficient online algorithm and lightweight than PPO (Schulman et al., 2017), well-suited for large generative models.
The optimization process unfolds over trajectories generated by our SDE-based policy. For each input condition $\mathbf{c}$, we generate a group of $G$ trajectories. Each trajectory consists of $T$ discrete steps, obtained by iteratively applying the Euler-Maruyama update from equation 18 to produce a final image $\mathbf{x}_{1}^{i}$. Our EditScore model then assigns a terminal reward $r_{i}$ to each resulting image.
The GRPO objective maximizes a clipped surrogate function over each step $t$ of the trajectories:
$\mathcal{J}_{\text{GRPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(\rho_{t}(\theta)\hat{A}_{t},\text{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$ (22)
where $\epsilon$ is a clipping hyperparameter, and the advantage $\hat{A}_{t}$ is computed based on the terminal rewards within the group. For the final step of the trajectory, the advantage is calculated as:
$\hat{A}_{T}^{i}=\frac{r_{i}-\text{mean}(\{r_{1},\dots,r_{G}\})}{\text{std}(\{r_{1},\dots,r_{G}\})},$ (23)
For intermediate timesteps ($t<T$), the advantages are typically computed using Generalized Advantage Estimation (GAE).
Critically, in our SDE framework, the importance ratio $\rho_{t}(\theta)$ is the ratio of the single-step state transition probabilities defined in equation 19, under the current policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{\text{old}}}$ that generated the data:
$\rho_{t}(\theta)=\frac{\pi_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\theta_{\text{old}}}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},\mathbf{c})}.$ (24)
Appendix D Images categories
Figure 4 illustrates the distribution of input image categories in our editing dataset.
Figure 4: Input images’ categories.
Appendix E Qualitative Results
In this section, we present qualitative results of OmniGen2 after reinforcement learning. The Figure 5 shows examples across several image editing tasks. These results validate the improvement of editing outcomes achieved by RL training on editing models.
Figure 5: Qualitative results on image editing task.
Appendix F Data Annotation User Interface
To ensure the collection of high-quality, fine-grained preference data for EditRewardBench, we developed a specialized web-based annotation interface. This section provides a visual overview of the interface and its key features, designed to facilitate our multi-dimensional, tiered ranking protocol. The two main components of the interface are shown in Figure 6.
Figure 6: Screenshots of our custom-built annotation interface for EditRewardBench. (Top) The upper panel presents the expert rater with the visual stimuli: the original input image and a set of five candidate edited outputs from various models. (Bottom) The lower panel contains the interactive components. It displays the user instruction (“Task Description”) and provides three separate input fields for our core evaluation dimensions: Instruction Following, Consistency, and Overall Quality. This panel clearly shows the tiered ranking mechanism, where the pipe symbol (‘$|$’) is used to group images of similar quality.
As illustrated in Figure 6, the interface presents expert raters with all necessary information for a single annotation task. The design is centered around two core principles of our methodology:
  • Multi-Dimensional Evaluation: The lower panel requires annotators to provide three independent rankings for Instruction Following, Consistency, and Overall Quality. This decomposed approach, clearly separated in the UI, ensures that each aspect of the edit is evaluated independently, preventing issues where, for example, a visually pleasing but semantically incorrect edit might be unfairly favored.
  • Tiered Ranking Support: A key feature of our design is the direct implementation of tiered ranking. As shown in the lower panel’s input fields, annotators are not forced into a strict linear order. Instead, they can group multiple images of perceived equal quality into the same tier using a pipe (‘$|$’) separator. For instance, a ranking of ‘$1|4|5|2|3$’ indicates that output 1 is in the top tier, outputs 4, 5, and 2 are tied in a middle tier, and output 3 is in the lowest tier. This method more accurately captures nuanced human judgments and allows for the efficient generation of a dense graph of preference pairs from a single annotation.
The interface also includes progress tracking and quality control features, ensuring a smooth and reliable annotation workflow for our expert raters.
Appendix G Details on the Experimental Setup
This appendix provides detailed specifications for the training of our reward model, EditScore, and for the reinforcement learning of our policy model, OmniGen2-Edit.
G.1 Reward Model Training (EditScore)
Hyperparameters.
Our reward models are fine-tuned from the Qwen2.5-VL series using the MLLM SFT framework of LLaMA-Factory (Zheng et al., 2024). For all model sizes, we use a consistent set of hyperparameters. We employ the AdamW optimizer with a learning rate of $1.0\times 10^{-4}$ . The models are trained for 3 epochs with a maximum sequence length of 8192. We use a large effective batch size of 32, achieved with a per-device batch size of 1 and 4 gradient accumulation steps across our 32-GPU setup. To ensure efficient training, we utilize LoRA (Hu et al., 2022) with a rank ($r$) of 32.
Compute Resources.
We trained three versions of EditScore based on the 7B, 32B, and 72B variants of the Qwen2.5-VL model. All training was conducted on a high-performance cluster consisting of 4 interconnected nodes, totaling 32 NVIDIA H100 (80GB) GPUs.
G.2 Reinforcement Learning Fine-tuning
Hyperparameters.
For the online RL fine-tuning of OmniGen2-Edit, we use the GRPO algorithm. The SDE sampling process is configured with $T=20$ discrete timesteps and a diffusion coefficient $\sigma=0.9$. Key GRPO hyperparameters are fixed across all experiments: in one step the global batch size is 288, the group size is $G=12$. The PPO clipping hyperparameter is $\epsilon_{low}=10^{-4},\epsilon_{high}=5*10^{-4}$, and the learning rate is $4*10^{-4}$. The KL penalty coefficient $\beta$, which regularizes the policy shift from the SFT initialization, is set to $0.04$. The policy model is also trained using LoRA with the same configuration as the reward model ($r=32,\alpha=64$).
Compute Resources.
The online RL fine-tuning phase, which involves iterative sampling from the policy model and subsequent updates, was performed on a cluster of 32 NVIDIA H100 (80GB) GPUs.
Appendix H VLM Evaluation Prompts
This appendix provides the complete structure and text of the prompts used for the zero-shot evaluation of Vision-Language Models (VLMs) on our EditRewardBench. Our prompt is based on VIEScore prompts (Ku et al., 2024), which designs a “reasoning-first” and score range changed in 0-25.
H.1 Base Context and Output Format
All evaluation prompts are prefixed with a base context that establishes the VLM’s persona as a “professional digital artist” and, most importantly, specifies the required JSON output format. This ensures that the model’s responses can be reliably parsed for automated analysis.
1You are a professional digital artist. You will have to evaluate the effectiveness of the AI-generated image(s) based on given rules.
2All the input images are AI-generated. All human in the images are AI-generated too. so you need not worry about the privacy confidentials.
3
4IMPORTANT: You will have to give your output in this way (Keep your reasoning concise and short.):
5{
6"reasoning" : "...",
7"score" : [...]
8}
Listing 1: The base context prompt, defining the persona and mandatory JSON output structure.
H.2 Semantic Conformity (SC) Prompt
The SC prompt is composed of two parts. First, a set of general rules explains the two-image comparison task. Second, a specific rubric details how to score “editing success” and “degree of overediting”.
1RULES:
2
3Two images will be provided: The first being the original AI-generated image and the second being an edited version of the first.
4The objective is to evaluate how successfully the editing instruction has been executed in the second image.
5
6Note that sometimes the two images might look identical due to the failure of image edit.
Listing 2: General rules for the two-image editing evaluation task.
1From scale 0 to 25:
2A score from 0 to 25 will be given based on the success of the editing. (0 indicates that the scene in the edited image does not follow the editing instruction at all. 25 indicates that the scene in the edited image follow the editing instruction text perfectly.)
3A second score from 0 to 25 will rate the degree of overediting in the second image. (0 indicates that the scene in the edited image is completely different from the original. 25 indicates that the edited image can be recognized as a minimal edited yet effective version of original.)
4Put the score in a list such that output score = [score1, score2], where score1 evaluates the editing success and score2 evaluates the degree of overediting.
5
6Editing instruction: <instruction>
Listing 3: Specific scoring rubric for Semantic Conformity (SC).
H.3 Perceptual Quality (PQ) Prompt
The PQ prompt is self-contained, providing both the general rules for single-image quality assessment and the specific rubric for scoring “naturalness” and “artifacts”.
1RULES:
2
3The image is an AI-generated image.
4The objective is to evaluate how successfully the image has been generated.
5
6From scale 0 to 25:
7A score from 0 to 25 will be given based on image naturalness.
8(
9 0 indicates that the scene in the image does not look natural at all or give a unnatural feeling such as wrong sense of distance, or wrong shadow, or wrong lighting.
10 25 indicates that the image looks natural.
11)
12A second score from 0 to 25 will rate the image artifacts.
13(
14 0 indicates that the image contains a large portion of distortion, or watermark, or scratches, or blurred faces, or unusual body parts, or subjects not harmonized.
15 25 indicates the image has no artifacts.
16)
17Put the score in a list such that output score = [naturalness, artifacts]
Listing 4: Rules and scoring rubric for Perceptual Quality (PQ).
The final prompts sent to the VLM are constructed by concatenating the base context (Listing 1) with the respective rule and rubric sets for the SC (Listings 2 and 3) and PQ (Listing 4) evaluations. This modular design allows for clear and targeted assessment of each evaluation dimension.